# Gaussian-Forest: Hierarchical-Hybrid 3D Gaussian Splatting for Compressed Scene Modeling

### v1

@@ Abstract
The field of novel-view synthesis has recently witnessed the emergence of 3D Gaussian Splatting, which represents scenes in a point-based manner and renders through rasterization. This methodology, in contrast to Radiance Fields that rely on ray tracing, demonstrates superior rendering quality and speed. However, the explicit and unstructured nature of 3D Gaussians poses a significant storage challenge, impeding its broader application. To address this challenge, we introduce the Gaussian-Forest modeling framework, which hierarchically represents a scene as a forest of hybrid 3D Gaussians. Each hybrid Gaussian retains its unique explicit attributes while sharing implicit ones with its sibling Gaussians, thus optimizing parameterization with significantly fewer variables. Moreover, adaptive growth and pruning strategies are designed, ensuring detailed representation in complex regions and a notable reduction in the number of required Gaussians. Extensive experiments demonstrate that Gaussian-Forest not only maintains comparable speed and quality but also achieves a compression rate surpassing 10 times, marking a significant advancement in efficient scene modeling. Codes are available at https://github.com/Xian-Bei/GaussianForest.
1 Introduction
Figure 1: Quantitative comparison across 13 real-world scenes from three datasets on rendering quality, model size, and rendering speed. The size of each point in the figure indicates the corresponding model size (in MB). Our Gaussian-Forest (GF) excels in adeptly balancing rendering speed and model size. Across all scenarios, GF achieves the highest speed-to-size ratio, surpassing all baselines by a large margin while ensuring high-fidelity rendering quality.
@@ 1 Introduction
Over the past few years, there has been rapid development in the field of 3D vision, marked by the emergence of the Radiance Field technique designed for 3D scene representation and novel view synthesis. This development has not only established a solid foundation but also acted as a significant catalyst for further advancements. As a pioneering effort, NeRF Mildenhall et al. (2020) represents 3D scenes implicitly using Multi-Layer Perceptrons (MLPs) and employs ray tracing for rendering, resulting in high visual quality. However, this approach comes with the drawback of unacceptably slow speeds for both training and inference. Subsequent research endeavors have explored various explicit or hybrid scene representations Liu et al. (2020); Sun et al. (2022) to enhance computational efficiency. Nonetheless, as these methods continue relying on ray tracing, which necessitates dense sampling across thousands of rays even in empty spaces, they encounter challenges in achieving real-time rendering rates. This challenge becomes more prominent when facing practical requirements such as high resolution, large-scale scenes, and consumer-grade devices.
@@
As a recent revolutionary development, 3D Gaussian Splatting (3DGS) has introduced an explicitly point-based approach for scene representation, pivoting from ray tracing to rasterization for both training and rendering processes. This innovative shift has resulted in state-of-the-art visual quality and comparable training efficiency while significantly boosting rendering speed. However, it comes with a primary constraint, which lies in its substantial storage requirements. Typically, it necessitates millions of Gaussians to represent a scene, resulting in a huge model size that even reaches thousands of megabytes. Such resource-intensive demands pose a significant obstacle to its practical application, particularly in scenarios with limited resources and bandwidth.
@@
In response to this practical challenge, we propose Gaussian-Forest for compressed 3D scene representation, which models each Gaussian with significantly fewer parameters by organizing hybrid Gaussians in a hierarchical forest structure, while concurrently controlling their overall number via adaptive growth and pruning. Our approach is motivated by the substantial parameter redundancy observed among the millions of Gaussians employed in 3DGS Kerbl et al. (2023), where groups of Gaussians exhibit implicit associations and share similar attributes. As demonstrated in Fig. 1, Gaussian-Forest adeptly balances storage, speed, and quality. The success of Gaussian-Forest hinges on three pivotal elements.
@@
Secondly, we organize hybrid Gaussians in a hierarchical manner, conceptualizing them as a forest for scene modeling. Explicit and implicit attributes of hybrid Gaussians are designated as leaf and non-leaf nodes, respectively, and interconnected through efficient pointers. In this formulation, through recursive tracing pointers upwards, each leaf node follows a unique path leading to the root of its corresponding tree, and all nodes along this path uniquely characterize a hybrid Gaussian. In addition, implicit attribute nodes at higher levels are significantly fewer in number (the quantity of root nodes is around 2% of the leaves) and are reused by a larger set of hybrid Gaussians, leading to a more compact scene representation while preserving adaptability and expressive capability.
@@
Thirdly, we propose an adaptive growth and pruning strategy for Gaussian-Forest. This dynamic growth relies on cumulative gradients to discern regions characterized by under-reconstruction or high uncertainty, such as object boundaries or regions with notable view-dependency. The expansion of new nodes in these complex regions facilitates swift scene adaptation, even with sparse or imprecise initial points. Simultaneously, regularly identifying and pruning insignificant leaves and branches, such as trivial Gaussians in simple regions like backgrounds, provides effective control over the total number of nodes. This ensures concise representations without compromising rendering quality, while contributing to the acceleration of both training and rendering. The primary contributions of this paper are summarized as follows:
@@
  • Introducing Gaussian-Forest, which represents a scene as a forest composed of hybrid Gaussians. By modeling each Gaussian with significantly fewer parameters, remarkable compactness is achieved while adaptability and expressiveness are retained.
@@
  • Developing adaptive growth and pruning strategies specifically tailored for Gaussian-Forest, facilitating rapid scene adaptation while avoiding unnecessary expansion of the number of Gaussians.
@@
  • Extensive experiments showcase Gaussian-Forest’s consistent attainment of comparable rendering quality and speed with a compression rate exceeding 10×10$\times$10 ×, strongly affirming its efficacy as an efficient technique for scene representation.
2 Related Work
Neural Radiance Field (NeRF) was introduced as a milestone for scene representation and novel view synthesis by Mildenhall et al. (2020). Subsequent work has since amplified this concept, with notable strides made across various fronts, such as regularization Kim et al. (2022); Niemeyer et al. (2022), supervision Wei et al. (2021); Deng et al. (2022); Yu et al. (2022), and extended applications like dynamics Pumarola et al. (2021); Cao and Johnson (2023) and editing Yuan et al. (2022); Zhou et al. (2023). Among these, the evolution in the scene representation and rendering paradigms stand as particularly profound advancements.
2.0.1 Implicit and Explicit Scene Representation
@@
NeRF Mildenhall et al. (2020) employs MLPs with compact size and continuous representation to model scenes. However, neural implicit representation leads to notably low time efficiency, attributed in part to long inference times of deep MLPs. Adhering to the principle of trading space for time, Plenoxels Wei et al. (2021) partitions the 3D space and stores associated attributes within each grid. However, high-resolution grids are necessary for detailed rendering, significantly escalating storage requirements. TensoRF Chen et al. (2022) applies the Tensor Decomposition on 3D grids, substantially reducing the model size but still remaining notably larger than implicit approaches. 3DGS Kerbl et al. (2023) represents a scene with collections of explicitly represented Gaussians, where millions of Gaussians are required for high-fidelity modeling, resulting in substantial model sizes.
2.0.2 Hybrid Scene Representation
@@
Typical hybrid representations entail the explicit storage of implicit features, which are inferred into concrete spatial attributes dynamically using neural networks. DVGO Sun et al. (2022) stores spatial features within volumetric 3D grids, albeit with huge space complexity. InstantNGP Müller et al. (2022) enhances this by incorporating multi-resolution 1D hash-tables, allowing positions with identical hashing values to share features. Point-NeRF Xu et al. (2022) adopts discrete 3D points to store spatial features. In summary, hybrid representations generally amalgamate the flexible nature of implicit representation and the high time efficiency of explicit ones. Similarly, our Gaussian-Forest is constructed based on a hybrid representation of 3D Gaussian.
2.0.3 Ray Tracing-based Rendering
@@
NeRF Mildenhall et al. (2020) trains and renders scenes via differentiable ray marching. Subsequent research has primarily adopted this rendering approach, where some endeavors have concentrated on enhancing rendering efficiency and geometry representation. InstantNGP Müller et al. (2022) maintains cascade occupancy grids to skip ray marching in empty space. Mip-NeRF360 Barron et al. (2022) introduces a proposal network to provide a rapid and approximate scene estimation. VolSDF Yariv et al. (2021) and NeuS Wang et al. (2021) extend this rendering to represent signed distance function (SDF) fields, achieving high-quality surface reconstruction. However, the dense sampling nature of ray tracing poses challenges in meeting real-time rendering demands.
@@
Figure 2: Overview of the proposed Gaussian-Forest. Gaussian-Forest hierarchically represents a scene as a forest composed of hybrid Gaussians, where non-leaf nodes capture their implicit attributes, while leaf nodes characterize explicit ones. Initiated from a compact set of singly linked lists, Gaussian-Forest adaptively grows in complex regions based on cumulative gradients to swiftly fit the scene. Leaf nodes with scaling and opacity below certain thresholds are considered trivial and subsequently removed. Such node count control ensures compact representations without compromising rendering quality while contributing to the acceleration of both training and rendering.
2.0.4 Rasterization-based Rendering
@@
Recent advancements have propelled differentiable rasterization Loper and Black (2014); Li et al. (2018); Yifan et al. (2019) into the forefront. In Kato et al. (2018), an approximate gradient solution for differentiable silhouette rasterization is introduced. Liu et al. (2019) formulates view rendering as an aggregation function, paving the way for learning full attributes of mesh from color supervisions. Diverging from these polygon mesh-based approaches, Pulsar Lassner and Zollhofer (2021) proposes a fast differentiable rasterizer based on 3D spheres. Inspired by this innovation, 3DGS Kerbl et al. (2023) further improves by employing anisotropic 3D Gaussians instead of isotropic spheres and a rasterizer that respects visibility ordering. Rasterization’s fast rendering speed and ease of integration with modern hardware position it as a promising avenue for further research.
3 Method
3.1 Preliminaries and Task Formulation
@@
3D Gaussian Splatting (3DGS) Kerbl et al. (2023) aims to model an arbitrary 3D scene with a set of anisotropic 3D Gaussians $\{\mathbf{G}_{i}\}_{i=1}^{N}$, which are initialized from structure-from-motion (SfM) sparse point clouds. Mathematically, each Gaussian is determined by the point mean position $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$ in 3D space:
@@
where $\mathbf{q}\in\mathbb{R}^{4},\mathbf{s}\in\mathbb{R}^{3}$ are covariance-related quaternion and scaling vectors. $\alpha\in\mathbb{R}$ stands for the opacity for the subsequent blending process. To account for color variations with viewing angles, each Gaussian’s color is modeled by 4-order spherical harmonics (SH), represented as $\mathbf{c}\in\mathbb{R}^{3\times 4^{2}}$. After training, the Gaussian parameters are determined, thus allowing the acquisition of the transformed 2D Gaussian on the image plane. Subsequently, a tile-based rasterizer is applied to sort $N$ Gaussians for $\alpha$-blending. Typically, modeling a real-world scene may require several million Gaussians. This substantial quantity, along with the 59 parameters associated with each Gaussian, significantly increases the storage requirements for the trained model and affects rendering speed during the sorting and blending process.
@@
To propose a more streamlined and practical solution, we propose Gaussian-Forest (GF), as illustrated in Fig. 2. Gaussian-Forest models Gaussian parameters within a hybrid tree, where each leaf node traces a distinct path to the root, thereby determining a specific Gaussian. Explicit attributes (e.g., position, opacity) are stored in the leaf nodes, while implicit attributes like scaling, rotation, and color are learned in the internal and root layers to maximize sharing across trees and reduce parameter redundancy (Sec. 3.2). Furthermore, to minimize the required number of Gaussians without compromising accuracy, the forests are dynamically grown and pruned (Sec. 3.3), focusing on adding siblings to accurately represent complex regions only.
3.2 Gaussian-Forest Modeling
@@
The objective of building the Gaussian-Forest structure is to enable individual Gaussians to retain their unique explicit properties for capturing distinct local areas, while simultaneously sharing common implicit attributes across the scene to reduce the number of free parameters represented in Eq. (2). Towards this, we realize this structure through a tree composed of $L$ layers, with $L$ set to 3 without loss of generality. Higher levels of trees will contain fewer nodes with features of higher dimensions to maximize the benefits of sharing. Herein, we introduce the hybrid modeling of tree nodes.

### v2

# GaussianForest: Hierarchical-Hybrid 3D Gaussian Splatting for Compressed Scene Modeling

@@ Abstract
The field of novel-view synthesis has recently witnessed the emergence of 3D Gaussian Splatting, which represents scenes in a point-based manner and renders through rasterization. This methodology, in contrast to Radiance Fields that rely on ray tracing, demonstrates superior rendering quality and speed. However, the explicit and unstructured nature of 3D Gaussians poses a significant storage challenge, impeding its broader application. To address this challenge, we introduce the GaussianForest modeling framework, which hierarchically represents a scene as a forest of hybrid 3D Gaussians. Each hybrid Gaussian retains its unique explicit attributes while sharing implicit ones with its sibling Gaussians, thus optimizing parameterization with significantly fewer variables. Moreover, adaptive tree growth and pruning strategies are designed, ensuring detailed representation in complex regions and a notable reduction in the number of required Gaussians. Extensive experiments demonstrate that GaussianForest not only maintains comparable speed and quality but also achieves a compression rate surpassing 10 times, marking a significant advancement in efficient scene modeling. Codes will be available at https://github.com/Xian-Bei/GaussianForest.
Index Terms:
Novel view synthesis, 3D reconstruction, model compression, neural rendering.
I Introduction
@@
Over the past few years, there has been rapid development in the field of 3D vision, marked by the emergence of the Radiance Field technique designed for 3D scene representation and novel view synthesis. This development has not only established a solid foundation but also acted as a significant catalyst for further advancements. As a pioneering effort, NeRF [1] represents 3D scenes implicitly using Multi-Layer Perceptrons (MLPs) and employs ray tracing for rendering, resulting in high visual quality. However, this approach comes with the drawback of unacceptably slow speeds for both training and inference. Subsequent research endeavors have explored various explicit or hybrid scene representations to enhance computational efficiency. Nonetheless, as these methods continue relying on ray tracing, which necessitates dense sampling across thousands of rays even in empty spaces, they encounter challenges in achieving real-time rendering rates. This challenge becomes more prominent when facing practical requirements such as high resolution, large-scale scenes, and consumer-grade devices.
Figure 1: Quantitative comparison across 13 real-world scenes from three datasets on rendering quality, model size, and rendering speed. The size of each point in the figure indicates the corresponding model size (in MB). Our GaussianForest (GF) excels in adeptly balancing rendering speed and model size. Across all scenarios, GF achieves the highest speed-to-size ratio, surpassing all baselines by a large margin while ensuring high-fidelity rendering quality.
@@
As a recent revolutionary development, 3D Gaussian Splatting (3DGS) [2] has introduced an explicitly point-based approach for scene representation, pivoting from ray tracing to rasterization for both training and rendering processes. This innovative shift has resulted in state-of-the-art visual quality and comparable training efficiency while significantly boosting rendering speed. However, it comes with a primary constraint, which lies in its substantial storage requirements. Typically, it necessitates millions of Gaussians to represent a scene, resulting in a huge model size that even reaches thousands of megabytes. Such resource-intensive demands pose a significant obstacle to its practical application, particularly in scenarios with limited resources and bandwidth.
@@
In response to this practical challenge, we propose GaussianForest for compressed 3D scene representation, which models each Gaussian with significantly fewer parameters by organizing hybrid Gaussians in a hierarchical forest structure, while concurrently controlling their overall number via adaptive growth and pruning. Our approach is motivated by the substantial parameter redundancy observed among the millions of Gaussians employed in 3DGS [2], where groups of Gaussians exhibit implicit associations and share similar attributes. As demonstrated in Fig. 1, GaussianForest adeptly balances storage, speed, and rendering quality. The success of GaussianForest hinges on three pivotal elements.
@@
Secondly, we organize hybrid Gaussians in a hierarchical manner, conceptualizing them as a forest structure for scene modeling. Explicit and implicit attributes of hybrid Gaussians are designated as leaf and non-leaf nodes, respectively, and interconnected through efficient pointers. In this formulation, through recursive tracing pointers upwards, each leaf node follows a unique path leading to the root of its corresponding tree, and all nodes along this path uniquely characterize a hybrid Gaussian. In addition, implicit attribute nodes at higher levels are significantly fewer in number (the quantity of root nodes is around 2% of that of the leaves) and are reused by a larger set of hybrid Gaussians, leading to a more compact scene representation while preserving adaptability and expressive capability.
@@
Thirdly, we propose an adaptive growth and pruning strategy for GaussianForest. This dynamic growth relies on cumulative gradients to discern regions characterized by under-reconstruction or high uncertainty, such as object boundaries or regions with notable view-dependency. The expansion of new nodes in these complex regions facilitates swift scene adaptation, even with sparse or imprecise initial points. Simultaneously, regularly identifying and pruning insignificant leaves and branches, such as trivial Gaussians in simple regions like backgrounds, provides effective control over the total number of nodes. This ensures concise representations without compromising rendering quality, while contributing to the acceleration of both training and rendering. The primary contributions of this paper are summarized as follows:
@@
  • Introducing GaussianForest, which represents a scene as a forest composed of hybrid Gaussians. By modeling each Gaussian with significantly fewer parameters, remarkable compactness is achieved while adaptability and expressiveness are retained.
@@
  • Developing adaptive growth and pruning strategies specifically tailored for GaussianForest, facilitating rapid scene adaptation while avoiding unnecessary expansion of the number of Gaussians.
@@
  • Extensive experiments showcase GaussianForest’s consistent attainment of comparable rendering quality and speed with a compression rate exceeding 10×10$\times$10 ×, strongly affirming its efficacy as an efficient technique for scene representation.
II Related Work
Neural Radiance Field (NeRF) was introduced by [1] as a milestone in scene representation and novel view synthesis. It utilizes neural networks to implicitly and continuously model 3D attributes and renders scenes through differentiable ray marching. Subsequent work has expanded upon this concept, making significant strides in various areas. For instance, [3, 4, 5] focus on enhancing the robustness of NeRF under adverse conditions, such as sparse view configurations and imperfect pose inputs. Meanwhile, [6, 7, 8, 9] incorporate depth information as priors to regularize the reconstruction of radiance fields. Furthermore, NeRF’s applications have been extended to dynamic scene reconstruction [10, 11, 12], scene editing [13, 14, 15, 16], and human head/face modeling [17, 18, 19]. Among all advancements, the evolution in the scene representation and rendering paradigms stand as particularly profound advancements. This section provides a brief review of the literature from these two perspectives.
II-A Scene Representation
II-A1 Implicit and Explicit Representation
@@
NeRF [1] utilizes MLPs characterized by compact size and continuous mapping to model scenes. This implicit representation has been embraced by numerous subsequent studies over time. However, in addition to the prolonged inference times associated with deep MLPs, updates at arbitrary positions require optimization across the entire network, further exacerbating its inefficiency. A straightforward acceleration strategy involves a trade-off between space and time: explicitly storing 3D attributes and retrieving them directly. Following this principle, Plenoxels [20] partitions the 3D space and stores the associated attributes within each grid. However, high-resolution grids are necessary for detailed rendering, which significantly increases storage requirements. TensoRF [21] addresses this issue by applying Tensor Decomposition to 3D grids, substantially reducing the model size, although it remains considerably larger than implicit approaches. Meanwhile, 3DGS [2] represents scenes using collections of explicitly represented Gaussians, where millions of Gaussians are required for high-fidelity modeling, resulting in substantial model sizes.
II-A2 Hybrid Representation
@@
Typical hybrid representations involve the explicit storage of implicit features, which are inferred into concrete spatial attributes using neural networks on-the-fly. For example, DVGO [22] stores spatial features within volumetric 3D grids, while Point-NeRF [23] uses discrete 3D points; both subsequently decode these features using MLPs. Generally, hybrid representations combine the flexible nature of implicit representations with the high time efficiency of explicit ones, but they still need to address the storage burden associated with explicit representation. InstantNGP [24] demonstrates an effective approach by incorporating multi-resolution 1D hash tables to enable feature sharing among positions with identical hashing values. NSVF [25] uses a sparse voxel octree to define voxel-bounded implicit fields for modeling local properties, enabling faster novel view rendering by skipping empty voxels. Similarly, our GaussianForest is constructed using a hybrid representation of 3D Gaussians and maximizes feature sharing by leveraging spatial redundancy to the fullest extent.
II-B Rendering Approach
II-B1 Ray Tracing-based Rendering
@@
NeRF [1] trains and renders scenes via differentiable ray marching, commonly known as volume rendering. While subsequent research primarily adopted this rendering approach, some work have made their endeavors on enhancing geometry fidelity [26, 27] and improving rendering efficiency [24, 28]. For instance, VolSDF [26] and NeuS [27] design the transformation between density and signed distance, extending volume rendering to represent SDF (signed distance function) fields for high-quality surface reconstruction. In pursuit of high efficiency, InstantNGP [24] maintains cascade occupancy grids to skip ray marching in empty space, and Mip-NeRF360 [28] introduces a proposal network to provide a rapid and approximate scene estimation. Nevertheless, the computationally intensive nature of dense sampling in ray tracing continues to challenge achieving real-time rendering capabilities.
@@
Figure 2: Illustration of the proposed GaussianForest. GaussianForest hierarchically represents a scene as a forest composed of hybrid Gaussians, where non-leaf nodes capture their implicit attributes, while leaf nodes characterize explicit ones. Initiated from a compact set of singly linked lists via K-Means, GaussianForest adaptively grows in complex regions based on cumulative gradients to swiftly fit the scene. Leaf nodes with scaling and opacity below certain thresholds are considered trivial and subsequently removed. Such node count control ensures compact representations without compromising rendering quality while contributing to the acceleration of both training and rendering.
II-B2 Rasterization-based Rendering
@@
Recent advancements have propelled differentiable rasterization [29, 30, 31, 32, 33, 34] into the forefront of computer graphics and computer vision. For instance, [29] introduces an approximate gradient solution for differentiable silhouette rasterization, enabling shape reconstruction from silhouette supervision. In [30], view rendering is formulated as an aggregation function that fuses the probabilistic contributions of all mesh triangles, facilitating the learning of full mesh attributes from color supervision. Diverging from these polygon mesh-based approaches, Pulsar [31] introduces a 3D sphere-based differentiable rasterizer that achieves unprecedented speed while avoiding topology problems. Inspired by Pulsar [31], 3DGS [2] further improves on the concept by employing anisotropic 3D Gaussians instead of isotropic spheres and a rasterizer that respects visibility ordering. The fast rendering speed and ease of integration with modern hardware make rasterization a promising avenue for further research. Our GaussianForest also follows this highly efficient rasterization approach.
III Method
III-A Preliminaries and Task Formulation
@@
3D Gaussian Splatting (3DGS) [2] aims to model a 3D scene with a set of $N$ anisotropic 3D Gaussians $\{\mathbf{G}_{i}\}_{i=1}^{N}$, which are initialized from structure-from-motion (SfM) sparse point clouds. Mathematically, each Gaussian is determined by the point mean position $\bm{\mu}$ and covariance matrix $\mathbf{\Sigma}$ in 3D space:
@@
where $\mathbf{q}\in\mathbb{R}^{4},\mathbf{s}\in\mathbb{R}^{3}$ are covariance-related quaternion and scaling vectors. $\alpha\in\mathbb{R}$ stands for the opacity for the subsequent blending process. To account for color variations with viewing angles, each Gaussian’s color is modeled by 4-order spherical harmonics (SH), represented as $\mathbf{c}\in\mathbb{R}^{3\times 4^{2}}$.
After training, the Gaussian parameters are determined, thus allowing the acquisition of the transformed 2D Gaussian on the image plane. Subsequently, a tile-based rasterizer is applied to sort $N$ Gaussians for $\alpha$-blending. Typically, modeling a real-world scene may require several million Gaussians. This substantial quantity, along with the 59 parameters associated with each Gaussian, significantly increases the storage requirements for the trained model and affects rendering speed during the sorting and blending process.
@@
To propose a more streamlined and practical solution, we propose GaussianForest, as illustrated in Fig. 2, which models Gaussian parameters within a hybrid tree, where each leaf node traces a distinct path to the root, thereby determining a specific 3D Gaussian. Explicit attributes including position and opacity are stored in the leaf nodes, while implicit attributes like covariance and color are learned in the internal and root layers to maximize sharing across trees and reduce parameter redundancy (Sec. III-B). Furthermore, to minimize the required number of Gaussians without compromising accuracy, the forests are dynamically grown and pruned (Sec. III-C).
III-B GaussianForest Modeling
@@
The objective of building the GaussianForest structure is to enable individual Gaussians to retain their unique explicit properties for capturing distinct local areas, while simultaneously sharing common implicit attributes across the scene to reduce the number of free parameters represented in Eq. (2). Towards this, we realize this structure through a tree composed of $L$ layers, with $L$ set to 3 without loss of generality. Higher levels of trees will contain fewer nodes with features of higher dimensions to maximize the benefits of sharing. Herein, we introduce the hybrid modeling of tree nodes.
III-B1 Hybrid Tree Nodes